# Progressive Multi-Source Domain Adaptation for Personalized Facial Expression Recognition

### v1

@@ V-E Impact of Approaches to Preserve Samples for Replay Domain
To evaluate the efficacy of our proposed approach, we conducted an extensive ablation study on the BioVid dataset, comparing various techniques to select relevant samples for the replay domain, illustrated in Fig. 3. We analyze performance across five different settings: No Preserve, Preserve Random Samples, Closest K-means, Closest DBSCAN (100/subject), and Ours, In No Preserve setting, we introduce a new source subject to the target without preserving any previous samples. At each introduction of the source subject, the model only relies on the current subject, which raises the forgetting issue and affects the adaptation process Preserve Random Samples, we randomly select and retain 2,000 samples from previous sources. This selection criterion performed the worst as compared to other techniques because the model fails to capture subject-specific features crucial for the target subject. This indicates that the model exhibits high sensitivity to person-specific characteristics. Thus, when presented with samples that deviate significantly from the target subject, the model struggles to adapt effectively. This highlights the importance of the sample selection strategy. Next, In Closest K-means, we employ the K-means clustering algorithm to identify and select the most relevant data points. Overall, the results from the previous technique are improved. However, in K-means, it is required to define the number of clusters, and its sensitivity to the outliers may limit its efficacy in selecting the most pertinent samples. Closest DBSCAN (100/subject), which utilizes the Density-Based Spatial Clustering of Applications with Noise (DBSCAN) algorithm to select 100 samples from each source subject. This approach demonstrates a significant improvement over K-means, primarily due to its ability to form clusters based on density measures without requiring predefined clusters while effectively eliminating outliers. This aligns with our objective of selecting samples closer to the centroid and avoiding outliers without prior knowledge of cluster numbers. Finally, our proposed method applies DBSCAN to create a density-based dictionary that stores the most relevant samples, eliminating the fixed per-subject sample selection. This highlights the importance of selecting the most expressive samples while avoiding the inclusion of potentially irrelevant samples from all source subjects.
@@ V-F Visualization
Density-based Selection of Replay Domain. Fig. 4 illustrates a t-SNE [69] representation of the embeddings from the last layer of the feature extractor ${F}(.)$ on Biovid target Subject-1 and Subject-5. We show the latent space of the features from a density-based selection of the replay domain before and after selecting a pertinent sample from a new source subject while adapting to a target subject. In the initial training step, the model retains more samples from the newly added source, as the network is more flexible in the initial stage, allowing more relevant samples to be included in a replay dictionary. On the other hand, as training progresses to step n, the replay dictionary is less affected by the newly added source subject. Two notable observations can be made here. First, the model gradually acquires useful knowledge from earlier subjects that are more closely related to the target, resulting in fewer samples being drawn from later sources, thereby minimizing the influence of more distant subjects. Second, the samples within the replay dictionary become tightly clustered with the target features, improving alignment with the target domain. Selection of Relevant Source Domains. Fig. 5 illustrates an example of four target subjects: Woman 27 (Sub-1), Man 36 (Sub-2), Woman 65 (Sub-9), and Man 25 (Sub-6) with their respective closest source subjects (using cosine similarity) that are structurally more similar. These subjects are selected first to optimize the model for target adaptation.
@@ V-G Complexity Analysis
Fig. 6 (c) presents the convergence time of the model to the target domain as the number of source domains increases incrementally. This experiment is conducted on the BioVid dataset, using 20 source domains and adapting to the target domain sub-4, with training performed for one epoch on an NVIDIA A100 GPU. We evaluate the time complexity across four approaches: All Previous Subjects, CMSDA, SImpAI, and P-MSDA (ours). In the All Previous Subjects method, all previously visited source domains are retained while progressively adding new ones, making it computationally expensive due to including all subjects at each step. In contrast, CMSDA and SImpAI integrate all source domains at the start of training. As these methods do not involve gradual adaptation, their time complexity is measured after processing all 20 source domains, resulting in a fixed time complexity regardless of the number of included domains. Our proposed approach, P-MSDA, maintains a constant number of three domains source, target, and replay throughout training. Gradual adaptation is achieved by replacing the current source with a new one and updating the replay domain with previously visited sources. During target adaptation, our method operates with $d$=$3$ domains and approximately $n$=$2000$ samples per domain, resulting in a total sample size $N=n\cdot d$. Training for one epoch takes approximately $t$=$15\cdot d$ seconds. In the All Previous Subjects method, each addition of a new source domain increases the total sample size to $N^{*}$=$n\cdot(d+1)$, leading to a proportional rise in training time to $t^{*}$=$t\cdot(d+1)$. However, our method ensures stable training complexity by consistently maintaining $d$=$3$ domains.
@@ VII-A Selection of Source Domains
Initialize: $P\leftarrow[]$#\## List to store domain distances
Initialize: $V\leftarrow[]$#\## List of adapted (visited) sources
Initialize $\mathbf{R}\leftarrow\emptyset$triangleright $\mathbf{R}$ is a replay domain for $(\mathbf{x},y)$ pairs
@@
4:          Compute cosine similarity $p_{a}$ in mini-batch between samples in $\mathbf{S}_{a}$ and $\mathbf{T}$ 3
@@
7:     Obtain $\widetilde{P}$ by normalizing $P$ between range $[0-1]$
@@
13:          ##\## Initially when $\mathbf{R}$ is $\emptyset$, assign $\mathbf{\widetilde{S}}_{a}$
@@ VII-B Density-based Selection of Replay Domain
Initialize: $E\leftarrow[]$#\## List to store distances
Initialize: $Z^{\text{s}}\leftarrow[]$, $Z^{\text{t}}\leftarrow[]$#\## List to store samples
@@ VIII Additional Implementation Detail
Our method is implemented on PyTorch [70]. For the training of source and target domain subjects, we selected a batch size of 16 with a momentum of 0.9. The learning rate was set to 0.0001, whereas the learning rate between linear and classification layers was set to 0.001. SGD optimizer is chosen with the weight decay of 5e-4. In the training of source subjects, the value of trade-off parameter $\lambda$ is set to 0.1. For the source selection, we set the threshold to 0.90. The selected source domains are used for the target adaptation, then the next batch of sources were calculated, and the adaptation process will continue until it reaches $top_{s}$ which is set to $40$ subjects, empirically see section V-F. Target Subjects: To make it consistent and comparable across all of our experiments. We select 10 fixed target subject domains from the BioVid Heat Pain Dataset [63]. To have a better target adaptability, we did not select any of the subjects that belong to the 20 unexpressed individuals as cited in [65]. These subjects are: 081014_w_27, 101609_m_36, 112009_w_43, 091809_w_43, 071309_w_21, 073114_m_25, 080314_w_25, 073109_w_28, 100909_w_65, 081609_w_40. For UNBC-McMaster we selected 5 target subjects: 107-hs107, 109-ib109, 121-vw121, 123-jh123, 115-jy115. For cross-dataset, we use same 10 target subjects from Biovid.
@@
TABLE IV: Different source selection criteria. N-Classes, determine by training a $N$ source classes; Maximum Mean Discrepancy (MMD), select closest sources from the target; Cosine-Similarity (CoS), pick based on the similarity score.
@@
TABLE V: Selecting $Top_{s}$ value for the selection of source domains
@@
IX Ablations
IX-A Impact of Selection of Source Domain
@@
In this ablation, we study different ways of selecting source subjects that are closer to the target domain in the BioVid dataset, shown in Table IV. In N-Classes, we train a ResNet18 model on the ’N’ number of classes, where ’N’ is the number of source subjects, i.e., 77. The model is trained for 20 epochs, where each class is associated with a source subject. After training, we introduce a target subject to make a prediction. The class with the highest prediction rate will be closest to the target subject. Therefore, we rank the subjects from the highest to the lowest prediction and adapt to the target subject sequentially. This method matches the performance of Sub-4 with other techniques, achieving an average accuracy of 0.83. The following criterion was estimating the distance of sources with the target by the maximum mean difference (MMD), improving performance by 0.4. Finally, when we apply cosine-similarity, it exceeds the other two methods by converging to every target subject, significantly improving the adaptation rate.
IX-B Choosing Top-s Value for Source Domain Selection
@@
The closest source selection criterion is based on the $\tau$, which is set to 0.80, and the $top_{s}$ is set to $40$ for Biovid subjects (except for sub-2). For UNBC-McMaster, the total number of source subjects is 20, selecting $top_{s}$ to 20 subjects. Table V shows empirically the selection criteria of $Top_{s}$ value. The experiment demonstrates the convergence of the proposed method, with 9 out of 10 target subjects achieving convergence within the first 40 source subjects ($Top_{s}=40$). Sub-2 displayed a delayed convergence, which required additional sources to reach stability. This outlier behavior indicates domain-specific characteristics present in the Sub-2 data distribution, requiring more sources of information for the convergence.
IX-C Visualization of Density-based Selection of Relevant Samples

### v2

@@ V-E Impact of Approaches to Preserve Samples for Replay Domain
To evaluate the efficacy of our proposed replay strategy, we conducted an ablation study on the BioVid dataset by comparing different techniques for selecting relevant samples, as shown in Fig. 4. We considered five settings: No Preserve, Preserve Random Samples, Closest K-means, Closest DBSCAN (100/subject), and Ours. In the No Preserve setting, new source subjects are introduced without retaining any previous samples, which results in severe forgetting. In Preserve Random Samples, 2,000 samples are randomly selected and retained; however, this performs poorly since random samples fail to capture subject-specific features crucial for adaptation. This highlights the need for a principled sample selection strategy. Next, in Closest K-means, we apply K-means clustering to identify representative samples. While this improves performance, K-means requires the number of clusters to be specified and is sensitive to outliers, limiting its effectiveness. In contrast, Closest DBSCAN (100/subject) uses the DBSCAN clustering algorithm, which does not require a predefined number of clusters and is more robust to outliers, as it relies instead on $\epsilon$ (neighborhood radius) and $minPts$ (minimum samples) to define density and guide cluster formation. This results in a marked improvement over K-means, supporting the use of density-based clustering. Finally, our proposed method extends this idea by applying DBSCAN to build a density-based dictionary of representative samples across subjects, eliminating fixed per-subject constraints and further improving adaptation.
@@ V-F Visualization
Density-based Selection of Replay Domain. Fig. 10 illustrates a t-SNE [72] representation of the embeddings from the last layer of the feature extractor F(.){F}(.) on Biovid target Subject-1 and Subject-5. We show the latent space of the features from a density-based selection of the replay domain before and after selecting a pertinent sample from a new source subject while adapting to a target subject. In the initial training step, the model retains more samples from the newly added source, as the network is more flexible in the initial stage, allowing more relevant samples to be included in a replay dictionary. On the other hand, as training progresses to step n, the replay dictionary is less affected by the newly added source subject. Two notable observations can be made here. First, the model gradually acquires useful knowledge from earlier subjects that are more closely related to the target, resulting in fewer samples being drawn from later sources, thereby minimizing the influence of more distant subjects. Second, the samples within the replay dictionary become tightly clustered with the target features, improving alignment with the target domain. Selection of Relevant Source Domains. Fig. 6 illustrates an example of four target subjects: Woman 27 (Sub-1), Man 36 (Sub-2), Woman 65 (Sub-9), and Man 25 (Sub-6) with their respective closest source subjects (using cosine similarity) that are structurally more similar. These subjects are selected first to optimize the model for target adaptation.
@@ V-G Complexity Analysis
[…]
V-H Limitation of P-MSDA
In the cross-dataset setting, certain target subjects (e.g., Sub-2, Sub-8, and Sub-10) showed lower adaptation performance compared to within-dataset adaptation. This underperformance is mainly due to demographic imbalances and missing expression classes in the source pool, which create a large domain shift between source and target subjects. As shown in Fig. 8, adapting from UNBC-McMaster (source) to BioVid (target) for Sub-2, a young Caucasian male, the first 4 selected sources were older Caucasian women and men. Although closest in learned feature space, these sources offered limited gains due to semantic mismatch. As progressive adaptation began, the model initially struggled to converge due to the large domain shift between source and target, which limited early performance gains. However, as training progressed, the framework gradually reduced this shift, leading to steady improvements over time. This highlights that while P-MSDA may converge more slowly when initial sources introduce a large domain gap, it can still recover as training progresses. Similar cases are observed in the in-the-wild datasets Aff-Wild2 and BAH, where class imbalance and large domain gaps further constrain adaptation. A detailed subject-wise and class-level analysis of these cases is provided in the supplementary material.
Figure 8: Adaptation behavior for Sub-2 (young Caucasian male) in cross-dataset. Early adaptation stages select demographically mismatched sources (older Caucasian women) with limited performance gain. In later stages, more structurally relevant sources (older Caucasian males) are selected, leading to a marked accuracy increase.
@@ VII-A Selection of Source Domains
To estimate the similarity matrix for all the source and target domain pairs and selecting the top-N samples,
@@
Initialize: $P\leftarrow[]$ #\# List to store domain distances
Initialize: $V\leftarrow[]$ #\# List of adapted (visited) sources
Initialize $\mathbf{R}\leftarrow\emptyset$ $\triangleright$ $\mathbf{R}$ is a replay domain for $(\mathbf{x},y)$ pairs
@@
4:    Compute cosine similarity $p_{a}$ in mini-batch between samples in $\mathbf{S}_{a}$ and $\mathbf{T}$ 14
@@
7:  Obtain $\widetilde{P}$ by normalizing $P$ between range [01][0-1]
@@
13:    #\# Initially when $\mathbf{R}$ is $\emptyset$, assign $\mathbf{\widetilde{S}}_{a}$
@@ VII-B Density-based Selection of Replay Domain
Estimate the Euclidean distance between each sample and centroid as,
Select the closest distance with the cluster centriods and create replay domain,
$\mathbf{R}^{*}=\{\mathbf{x}^{\text{r}}_{i},y^{\text{r}}_{i}\}^{N^{\text{r}}}_{i=1}$ (18)
@@
Initialize: $E\leftarrow[]$ #\# List to store distances
Initialize: $Z^{\text{s}}\leftarrow[]$, $Z^{\text{t}}\leftarrow[]$ #\# List to store samples
@@ VIII Additional Implementation Detail
Our method is implemented on PyTorch [73]. For the training of source and target domain subjects, we selected a batch size of 16 with a momentum of 0.9. The learning rate was set to 0.0001, whereas the learning rate between linear and classification layers was set to 0.001. SGD optimizer is chosen with the weight decay of 5e-4. In the training of source subjects, the value of trade-off parameter $\lambda$ is set to 0.1. For the source selection, we set the threshold to 0.90. The selected source domains are used for the target adaptation, then the next batch of sources were calculated, and the adaptation process will continue until it reaches $top_{s}$ which is set to $40$ subjects, empirically see section X-E. Target Subjects: To make it consistent and comparable across all of our experiments. We select 10 fixed target subject domains from the BioVid Heat Pain Dataset [64]. To have a better target adaptability, we did not select any of the subjects that belong to the 20 unexpressed individuals as cited in [68]. These subjects are: 081014_w_27, 101609_m_36, 112009_w_43, 091809_w_43, 071309_w_21, 073114_m_25, 080314_w_25, 073109_w_28, 100909_w_65, 081609_w_40. For UNBC-McMaster we selected 5 target subjects: 107-hs107, 109-ib109, 121-vw121, 123-jh123, 115-jy115. For cross-dataset, we use same 10 target subjects from Biovid.
VIII-A Implementation Detail
In all experiments, we employ the ResNet18 backbone [71], which consists of the encoder $F$ and the discriminative component $C$. To adapt $F$ for subject-based MSDA, we follow the same protocol as  [18] to remove the first ReLU, the MaxPool layers, and the final 2D adaptive average pooling layer. In our experiments, the backbone is shared across every domain, followed by the shared classifier. The images are resized to $\times100$ resolution, and the model is trained with stochastic gradient descent (SGD) with a batch size of 16 and a learning rate of $10^{-4}$. For generating target pseudo-labels we set a threshold based on $\theta=\theta_{0}-\delta\left\lfloor\frac{e}{N}\right\rfloor$, the initial value of $\theta_{0}$ is set to 0.91, that was updated after every $N=20$, with the reduction value $\delta$ is set to 0.01. For the ACPL technique to generate reliable target PLs, we follow the same setting as Zeeshan et al. [18] and use a horizontal flip as an augmented version of the image. For replay relevant samples, we set $top_{n}$ and $N_{r}$ to 2000 samples, which are updated after each newly added source subject.
TABLE IV: Aff-Wild2 accuracy on 10 target subjects for different baselines and our method.
Source-only 0.12 0.16 0.14 0.25 0.17 0.74 0.29 0.31 0.13 0.53 0.28
CMSDA [20]
Sub-based [18]
Sub-basedtop-k [18]
0.21
0.11
0.12
0.16
0.17
0.18
0.18
0.20
0.16
0.20
0.32
0.24
0.20
0.26
0.20
0.56
0.85
0.88
0.31
0.27
0.32
0.52
0.55
0.76
0.21
0.15
0.13
0.37
0.62
0.65
0.29
0.35
0.36
P-MSDA (Ours) 0.14 0.29 0.14 0.57 0.49 0.78 0.66 0.63 0.12 0.76 0.46
TABLE V: BAH accuracy on 10 target subjects with baselines and our method.
Source-only 0.34 0.59 0.32 0.71 0.67 0.37 0.44 0.51 0.70 0.47 0.51
Sub-based [18]
Sub-basedtop-k [18]
0.38
0.38
0.60
0.60
0.35
0.49
0.75
0.76
0.77
0.78
0.38
0.46
0.47
0.45
0.53
0.51
0.71
0.74
0.50
0.50
0.54
0.57
P-MSDA (Ours) 0.67 0.61 0.90 0.89 0.86 0.77 0.64 0.54 0.71 0.55 0.71
TABLE VI: Comparison between ResNet18 and ViT backbones on BioVid dataset.
 Backbone Methods Sub-1 Sub-2 Sub-3 Sub-4 Sub-5 Sub-6 Sub-7 Sub-8 Sub-9 Sub-10 Avg
ResNet18 Source-only 0.62 0.61 0.65 0.55 0.51 0.71 0.70 0.52 0.54 0.55 0.59
Sub-based [18] 0.93 0.69 0.84 0.66 0.60 0.76 0.84 0.55 0.62 0.66 0.71
Sub-basedtop-k [18] 0.93 0.71 0.86 0.87 0.88 0.92 0.86 0.77 0.84 0.68 0.83
P-MSDA 0.99 0.76 0.87 0.92 0.89 0.94 0.87 0.81 0.98 0.78 0.88
ViT Source-only 0.94 0.73 0.65 0.63 0.74 0.62 0.86 0.50 0.54 0.52 0.68
Sub-based [18] 0.94 0.74 0.60 0.88 0.83 0.69 0.83 0.52 0.58 0.53 0.71
Sub-basedtop-k [18] 0.93 0.70 0.63 0.85 0.81 0.68 0.86 0.51 0.59 0.53 0.71
P-MSDA 0.95 0.76 0.91 0.92 0.79 0.91 0.87 0.78 0.99 0.72 0.86
TABLE VII: Comparison between ResNet18 and ViT backbones on UNBC-McMaster dataset.
 Backbone Methods Sub-1 Sub-2 Sub-3 Sub-4 Sub-5 Avg
ResNet18 Source-only 0.74 0.84 0.81 0.68 0.83 0.78
Sub-based [18] 0.81 0.91 0.94 0.72 0.92 0.86
P-MSDA 0.87 0.93 0.94 0.74 0.94 0.88
ViT Source-only 0.79 0.89 0.47 0.73 0.93 0.76
Sub-based [18] 0.78 0.88 0.80 0.70 0.91 0.81
P-MSDA 0.82 0.93 0.90 0.71 0.94 0.86
@@ VIII-A Implementation Detail
TABLE VIII: Different source selection criteria. N-Classes, determine by training a $N$ source classes; Maximum Mean Discrepancy (MMD), select closest sources from the target; Cosine-Similarity (CoS), pick based on the similarity score.
@@
TABLE IX: Selecting $Top_{s}$ value for the selection of source domains
@@
IX Aditional Results
IX-A Results on Additional Datasets
We broadened our evaluation beyond pain-related datasets by including two additional datasets from different application domains: (1) Aff-Wild [66], a widely used in-the-wild emotion recognition dataset, and (2) the Behavior Ambivalence Hesitancy (BAH) dataset [67], which captures more naturalistic behaviors under partially uncontrolled conditions. For Aff-Wild, we treated each video as an independent “subject,” excluding any videos containing multiple individuals or duplicate identities to better approximate a subject-specific setting. A total of 10 videos were selected as target subjects, each containing a unique individual. To ensure a diverse emotion distribution, we hand-picked videos with the maximum number of distinct emotion classes (Table X). The remaining videos served as source subjects. Despite the absence of explicit subject identifiers and the inherent intra-subject variability, P-MSDA achieved an average accuracy of 0.46 (Table IV), outperforming all baselines and demonstrating strong adaptability under challenging in-the-wild conditions.
For BAH, which provides multiple recordings per subject in semi-uncontrolled environments. We selected 10 target subjects (Table XIII): 82711, 82687, 82585, 82592, 82598, 82632, 82681, 82683, 82708, 82714. The remaining subjects were treated as sources.
P-MSDA achieved an average performance of 0.71 (Table V), substantially higher than competing methods. These results demonstrate that our approach generalizes to both emotion-oriented FER and behavioral analysis scenarios, and that datasets with explicit subject information allow P-MSDA to fully leverage its personalization strengths. These additional experiments broaden our evaluation beyond pain datasets, demonstrating that P-MSDA generalizes to emotion-oriented FER scenarios and in-the-wild conditions, while also highlighting that datasets with explicit subject identifiers and multiple recordings per subject remain better suited to fully leveraging its personalization capabilities.
TABLE X: Class distribution across 10 target subjects of Aff-Wild2. Sub-9 (highlighted) shows a pronounced imbalance.
 Subject Video-ID #Classes Neutral Anger Disgust Fear Happiness Sadness Surprise Total
Sub-1 9-15-1920x1080 7 3483 631 362 1535 3019 22615 2272 33917
Sub-2 7-60-1920x1080 7 1914 1233 2098 3630 4153 144 1027 14199
Sub-3 video73 7 139 173 140 127 1686 156 3355 5776
Sub-4 8-30-1280x720 6 2204 1798 716 30 894 0 1369 7011
Sub-5 198 6 192 48 121 287 323 0 825 1796
Sub-6 384 6 1191 0 27 55 57 20 683 2033
Sub-7 video34 6 2135 16 99 0 5895 535 102 8782
Sub-8 147 5 $30$ 0 0 16 81 1600 16 4743
Sub-9 121-24-1920x1080 5 918 206 0 5546 281 0 414 7365
Sub-10 131 5 542 72 62 0 2362 0 39 3077
IX-B Comparison between ResNet and ViT Backbones
We extended our evaluation to include a transformer-based backbone, specifically the Vision Transformer (ViT), alongside the original ResNet18. On the BioVid dataset Table VI, the source-only ViT achieves an average accuracy of 0.68, outperforming ResNet18 0.59, shows that transformers can provide stronger baselines under limited adaptation. However, our proposed P-MSDA substantially boosts ResNet18 to 0.88 and ViT to 0.86, with consistent improvements across subjects, confirming that the method enhances both CNNs and transformer based models. On the UNBC-McMaster dataset Table VII, ResNet18 begins with a stronger source-only baseline (0.78) compared to ViT (0.76), but again P-MSDA improves both backbones, achieving 0.88 with ResNet18 and 0.86 with ViT. These results demonstrate that while ViT can offer competitive baselines, the performance gains from P-MSDA are not dependent on a specific backbone, highlighting that the method is backbone-agnostic and effective for both convolutional and transformer architectures.
X Ablations
X-A Impact of Selection of Source Domain
@@
In this ablation, we study different ways of selecting source subjects that are closer to the target domain in the BioVid dataset, shown in Table VIII. In N-Classes, we train a ResNet18 model on the ’N’ number of classes, where ’N’ is the number of source subjects, i.e., 77. The model is trained for 20 epochs, where each class is associated with a source subject. After training, we introduce a target subject to make a prediction. The class with the highest prediction rate will be closest to the target subject. Therefore, we rank the subjects from the highest to the lowest prediction and adapt to the target subject sequentially. This method matches the performance of Sub-4 with other techniques, achieving an average accuracy of 0.83. The following criterion was estimating the distance of sources with the target by the maximum mean difference (MMD), improving performance by 0.4. Finally, when we apply cosine similarity, it exceeds the other two methods by converging to every target subject, significantly improving the adaptation rate.
X-B Choosing Top-s Value for Source Domain Selection
@@
The closest source selection criterion is based on the $\tau$, which is set to 0.80, and the $top_{s}$ is set to $40$ for Biovid subjects (except for sub-2). For UNBC-McMaster, the total number of source subjects is 20, selecting $top_{s}$ to 20 subjects. Table IX shows empirically the selection criteria of $Top_{s}$ value. The experiment demonstrates the convergence of the proposed method, with 9 out of 10 target subjects achieving convergence within the first 40 source subjects ($Top_{s}=40$). Sub-2 displayed a delayed convergence, which required additional sources to reach stability. This outlier behavior indicates domain-specific characteristics present in the Sub-2 data distribution, requiring more sources of information for the convergence.
TABLE XI: Effect of varying similarity threshold on accuracy and number of similarity recalculations.
Threshold ($\gamma$) Accuracy Re-calc.
0.50 0.93 0
0.60 0.95 1
0.70 0.97 2
0.80 0.99 3
0.90 0.95 8
1.00 0.94 40
TABLE XII: Effect of varying pseudo-label threshold ($\tau_{\text{PL}}$) on accuracy.
Threshold ($\tau_{0}$) Accuracy
0.50 0.87
0.60 0.93
0.70 0.95
0.80 0.92
0.90 0.99
1.00 0.94
Figure 11: BAH source demographic statistics for male, 25-34 age group.
TABLE XIII: Target subjects of the BAH dataset with demographic and class distributions. N: Neutral, A/H: Ambivalence/Hesitancy.
 Subject ID Sex Age Group Ethnicity Frames A/H Frames Total
Sub-1 82711 Male 25–34 South Asian 2160 4376 2160
Sub-2 82687 Female 25–34 South Asian 1550 1081 1550
Sub-3 82585 Male 25–34 Black 2895 406 2895
Sub-4 82592 Female 25–34 Black 3685 462 3685
Sub-5 82598 Female 25–34 White 5995 946 5995
Sub-6 82632 Female 45–54 White 4222 1263 4222
Sub-7 82681 Female 25–34 Middle Eastern 1734 1068 1734
Sub-8 82683 Female 35–44 East Asian 2165 2337 2165
Sub-9 82708 Male 25–34 East Asian 1229 536 1229
Sub-10 82714 Male 65+ White 1153 1169 1153
X-C Ablation on Similarity Threshold ($\gamma$).
For subject selection, we set the similarity threshold ($\gamma$) to $0.80$ in our main experiments. Initially, we select source subjects whose similarity score with the target subject exceeds this threshold. After adapting to the selected sources, we re-calculate similarity scores between the target and all remaining sources until a total of $top_{s}=40$ sources is reached, thereby progressively incorporating more closely aligned subjects. Note that an additional ablation on the choice of $N$ sources is provided in Section IV-B of the supplementary material. We conducted a study as shown in Table XII to determine the optimal value of the similarity threshold. If $\gamma$ is set too low, many sources are incorporated at once, which eliminates the opportunity to re-calculate similarities and progressively refine source selection. Conversely, if $\gamma$ is set too high, the system may re-calculate excessively, introducing noise from marginally similar sources. Setting $\gamma=0.80$ provides the best balance, yielding superior convergence and accuracy by controlling the number of re-calculations while progressively incorporating the most relevant sources.
X-D Ablation on Pseudo-Label (PL) Threshold ($\tau$).
Following prior MSDA literature [38, 20, 18], we chose a relatively high PL threshold to ensure reliable pseudo-label selection and minimize the effect of noisy labels. Moreover, we adopt an adaptive threshold $\tau=\tau_{0}-\delta\left\lfloor\frac{e}{U}\right\rfloor$ to select the value of confidence PL threshold. Specifically, $\tau$ is initialized with $\tau_{0}=0.90$ and is gradually decreased in a step-wise manner, where $e$ denotes the current epoch, $U=20$ is the update interval, and $\delta$ controls the decrement rate set to $0.01$. This schedule allows the model to start with highly confident pseudo-labels and progressively incorporate more samples as adaptation proceeds. We further performed an ablation on the initial $\tau_{0}$ value. As shown in Table XII, lower values (e.g., $\tau_{0}\leq 0.80$) significantly degrade performance due to the early inclusion of noisy pseudo-labels before the model is sufficiently adapted, leading to poor convergence. In contrast, a higher starting value (e.g., $\tau_{0}=0.90$) provides the best balance, enabling gradual adaptation while avoiding noise in the early stages.
X-E Visualization of Density-based Selection of Relevant Samples